This seemed simple enough but I keep screwing it up. I've had good help from this site and would really appreciate it again.
"Dana buys a peice of carpet that measures 20 sq. yds.. Will she be able to completely cover a rectengular floor that measures 12.5 ft. by 16.5 ft?"

Found 2 solutions by edjones, scott8148:
Answer by edjones(8007)

12.5*16.5=206.25 ft^2
1 yd^2=9 ft^2
206.25/9=22.9 yd^2 which is larger than the area of the carpet.
No. (answer)
.
Ed

Answer by scott8148(6628)

1yd=3ft __ squaring __ 1yd^2=9ft^2

so 20yd^2 is 180ft^2

12.5*16.5=206.25

not enough carpet